OpenStudy (magbak):

PLEASE HELP ME Medal will be awarded y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3? 1 9 27 81 -------------------------------------------------------------------------------- At a specialty nut store, almonds cost $4.30 per pound and walnuts are priced at $3.75 per pound. How many pounds of walnuts are in a 12-pound mixture priced at $47.75? 4 5 6 7 What is the fourth term in the expansion of (x – 4)6? –1280x3 –256x4 256x4 1280x3

OpenStudy (calculusxy):

The second one is a Guess, Check, and Revise problem. You have to think of two addends that will add to 12 pounds. So at first I tried out how much will the price be if there are 6 pounds of almonds and 6 pounds of walnuts. But then I figured out that the price was too much than required. Then I tried decreasing the amount of almond by one pound (which is 5 pounds) and keeping the amount of walnuts the same. So what I did is ($4.30x5)+($3.75x6)=$47.75.

OpenStudy (whpalmer4):

direct variation of thing 1 with thing 2 means that thing1 = k * thing 2, where k is a constant. so, if y varies directly with the square of x, that means \(y = kx^2\) Plug in the known data to find \(k\) and update the equation to find the other point. \[3=k(1)^2\]\[3=k\] \[y = 3x^2\]\[y = (3)^2 = 9\]

OpenStudy (calculusxy):

Sorry about the first question. I figured out that y has to be squared. So since y=9, the answer will be 81 because 9x9=81.

OpenStudy (whpalmer4):

You can guess and check for the second problem, or write the equations and solve. A is amount of almonds W is amount of walnuts A+W = 12 4.3A + 3.75W = 47.75 Solve by substitution or elimination. To substitute, solve the first equation for one of the variables in terms of the other, then plug into the second and solve: A+W = 12 A = 12-W 4.3(12-W) +3.75W = 47.75 51.6 - 4.3W + 3.75W = 47.75 W = 7 A = 12-W = 12-7 = 5 To eliminate, multiply each equation by a number that will cause one of the coefficients to be equal in magnitude, then add or subtract as appropriate. A + W = 12 4.3A + 3.75 W = 47.75 multiply first eq by -4.3 -4.3A -4.3W = -51.6 4.3 A + 3.75 W = 47.75 ------------------- -0.55 W = -3.85 W = 7

OpenStudy (whpalmer4):

For the 3rd problem, I would use the binomial theorem.

OpenStudy (calculusxy):

But isn't this a question related to variables?

OpenStudy (calculusxy):

I meant for the second question.

OpenStudy (whpalmer4):

what are A and W if not variables? well, okay, a fast food joint :-)

OpenStudy (whpalmer4):

@calculusxy "So since y=9, the answer will be 81 because 9x9=81." where are you getting that y = 9?

OpenStudy (magbak):

So what is number 3

OpenStudy (whpalmer4):

@magbak do you know the binomial theorem?

OpenStudy (magbak):

Also is number 1 p

OpenStudy (magbak):

IS number 1 y=9

OpenStudy (whpalmer4):

sorry, I missed a 3 on my answer to #1 we found k = 3, so the equation is \(y = 3x^2\) and we want to find \(y\) when \(x=3\) so that gives us \(y=3(3)^2 = 3*9 = 27\)

OpenStudy (magbak):

What is number 3 please.

OpenStudy (whpalmer4):

you didn't answer my question about the binomial theorem...do you know it?

OpenStudy (magbak):

Yes.

OpenStudy (whpalmer4):

okay, that's what I would use to find the answer. What is the fourth term in the expansion of \((x – 4)^6\)? We have \(a = x, b = -4, n = 6\) and we want to find the 4th term

OpenStudy (magbak):

Yes. But I believe pascals triangle will be the easiest.

OpenStudy (whpalmer4):

okay, the race is on :-)

OpenStudy (magbak):

OK

OpenStudy (magbak):

-6144x that is the 6th term.

OpenStudy (whpalmer4):

but we want the 4th term...

OpenStudy (whpalmer4):

you are correct about the 6th term, however I am unable to award extra credit :-)

OpenStudy (magbak):

I mean -1280x^3

OpenStudy (whpalmer4):

Our answer is going to be \[\binom{6}{3}a^{3}b^{6-3} = \frac{6!}{3!3!} (x)^3(-4)^3 = \frac{6*5*4*3*2*1}{3*2*1*3*2*1}x^3(-64) \]\[= 20(-64)x^3 = -1280x^3\]

OpenStudy (magbak):

Is the fourth trem

OpenStudy (whpalmer4):

agreed.

OpenStudy (magbak):

Yep thank you. Got to go. Thank you so much.

OpenStudy (whpalmer4):

you're welcome!